10 of the 6 years old kids were selected from a school and the height of each individual were shown as in the table




Individual Height (cm)



A 120.6



B 116.4



C 117.2



D 118.1



E 114.1



G 116.9



H 113.3



I 121.1



J 116.9



K 117.0




According to the national statistics report, the average height of 6 years old kids in the country is 120.0 cm. What is the T score of this sample?

1
Expert's answer
2021-11-29T12:59:07-0500
"mean=\bar{x}=\dfrac{1}{10}(120.6+116.4+117.2"

"+118.1+114.1+116.9+113.3"

"+121.1+116.9+117.0)=117.16"

"s^2=\dfrac{\sum_i(x_i-\bar{x})^2}{n-1}"

"=\dfrac{1}{10-1}((120.6-117.16)^2+(116.4-117.16)^2"

"+(117.2-117.16)^2+(118.1-117.16)^2"


"+(114.1-117.16)^2+(116.9-117.16)^2"

"+(113.3-117.16)^2+(121.1-117.16)^2"

"+(116.9-117.16)^2+(117.0-117.16)^2)"

"=\dfrac{53.244}{9}=5.916"

"t=\dfrac{\bar{x}-\mu}{s/\sqrt{n}}=\dfrac{117.16-120.0}{\sqrt{5.916}/\sqrt{10}}"

"\approx-3.692362"